### v2

@@ Abstract
{IEEEkeywords}
Affine frequency division multiplexing (AFDM), orthogonal frequency division multiplexing (OFDM), evolutionary, 6G and beyond (6G+), backward compatibility, system reusability.
[…]
1 Introduction
1.1 Background
@@ 1.2 Why AFDM?
1.3 Contributions of this article
@@
1.4 Organization of this article
The remainder of this article is structured as follows: Section 2 introduces the system and channel models adopted throughout the article, including the novel formulation of a general FDFD channel input-output model based on physical pulse footprints to capture inter-sample coupling effects in practical transceiver implementations accurately; Section 3 details the transceiver architecture of AFDM, highlighting its compatibility and reusability with respect to OFDM, considering various aspects, such as prefixing, pulse shaping, and resource management; Section 4 presents a variety of thorough analyses of AFDM in comparison with OFDM, including implementation complexity and scalability, channel estimation and detection architecture, robustness to impairments like phase noise (PHN) and CFO, and advanced functionalities including multiple-input multiple-output (MIMO), chirp-domain modulation and physical-layer security; Finally, Section 5 qualitatively summarizes the reusibility of the presented AFDM considerations and discussion across the RF, PHY, and higher layers, followed by concluding remarks in Section 6.
@@ 2 System Model
2.1 Continuous-Time SISO Doubly Dispersive Channel
@@
2.2 Discrete-Time Equivalent Model
@@
2.3 Generalized Matrix-form Input-Output Relationship
@@
where $\phi_{\mathrm{cp}}(\cdot)$ is the waveform-dependent phase-offset function of the prefix samples, elaborated in Section 3.
@@
2.4 Extension to MIMO Channels
@@
Following the sampling procedure in Section 2.2 and the FDFD case, the discrete-time baseband channel between the $n_{\mathrm{t}}$-th transmit and $n_{\mathrm{r}}$-th receive antennas can be written as
@@
3 On the Compatibility of AFDM over OFDM
@@
To that end, let us first introduce the input-output system model of the conventional OFDM and the AFDM modulator/demodulator, given the model of the discrete-time doubly dispersive channel in Section 2.2.
3.1 OFDM Signal Model
@@
Figure 3: Effective channel matrices $\mathbf{\Xi}$ for OFDM and AFDM, respectively on the first and second rows, in a 3-path doubly dispersive scenario with $N=64$. The physical channel parameters are defined by path coefficients $\{1,0.9,0.8\}$, normalized delays $\{1.3,3.25,5.96\}$, and normalized digital Dopplers $\{1.1,-2.3,0.85\}$, where for the IDID and IDFD cases, the delays and Dopplers are rounded to the nearest integers as relevant. For the AFDM implementation, chirp parameters are set to $c_{1}=(2(f_{\max}+1)+1)/2N$, with $f_{\max}=3$, and $c_{2}=1/2N$. It can be seen that on top of the discussed effects of the fraction taps in Section 2-2.3, the different paths of the OFDM effective channel matrices are indistinguishable under Doppler shifts, leading to significant inter-carrier interference, while AFDM maintains a more distinguishable effective channel matrix, preserving better orthogonality.
@@
3.2 AFDM Signal Model
@@
3.3 MIMO Extension of OFDM and AFDM Signal Models
@@
3.4 A Note on OFDM and AFDM Compatibility
@@
3.5 Pulse-Shaping, Windowing, and Prefixing Blocks
@@
4 Analysis of the AFDM Transceiver over OFDM
4.1 Modulation Complexity and Scalability
@@
Table 1: Computational complexity of key operations and total modulator/demodulator complexities in FLOPs, with $N=KL$.
@@
Comparing the modulator complexities including the pulse-shaping overheads of OFDM, AFDM, and the two OTFS implementations in (42) and (46), which are summarized in Table 1, it is evident that both OTFS approaches introduce significantly higher computational burdens compared to the OFDM due to their multi-stage transform structures and 2D processing requirements under practical systems [^3] Regarding FFT dimension flexibility, OFDM and AFDM support arbitrary FFT/IFFT sizes $N$ with one-dimensional frequency-domain processing, enabling flexible numerology and resource allocation. In contrast, OTFS imposes strict integer divisibility constraints on the delay-Doppler grid dimensions $K,L\in\mathbb{N}$ such that $N=KL$, limiting design flexibility and potentially resulting in awkward FFT sizes for practical implementations and efficient computations., while AFDM maintains a modest overhead over OFDM and therefore stands out as a more computationally efficient alternative.
@@
4.2 Pilot Scheme and Channel Estimation
@@
4.2.1 Pilot Schemes
Due to the structure of the effective channel matrix $\mathbf{\Xi}^{\mathrm{AFDM}}$ as described in Section 3.2, the usual pilot schemes for AFDM involve one or multiple pilot subcarriers [79, 29], each of which is isolated from other subcarriers by guard bands (in the DAFT domain) to capture the contribution of each path without interference properly.
@@ 4.2.1 Pilot Schemes
4.2.2 Channel Estimation in the DAFT Domain
@@
$\mathbf{r}=\bigg(\sum_{p=1}^{P}\sum_{q=-B}^{B}h_{p}\cdot\mathbf{\Phi}_{p}\cdot\mathbf{V}^{f_{p}}\cdot g_{q}\cdot\mathbf{\Pi}^{\ell_{p}+q}\bigg)\cdot\mathbf{s}+\mathbf{w}.$ (55)
@@
$\mathbf{r}=\Bigg(\sum_{u=1}^{P(2B+1)}\!\!\!h_{u}^{\prime}\cdot\mathbf{\Phi}_{u}\cdot\mathbf{V}^{f_{u}}\cdot\mathbf{\Pi}^{\ell_{u}^{\prime}}\Bigg)\cdot\mathbf{s}+\mathbf{w},$ (56)
@@
4.2.3 Special Case: OFDM-like Estimation Under Low Doppler
@@
4.3 Signal Detection and Receiver Architectures
@@
4.3.1 One-shot Receivers (No outer iterations)
@@
4.3.2 Turbo Receiver
Alternative to the one-shot receivers where there are no outer iterations between the detector and the decoder, AFDM turbo receivers have been investigated in [100, 95, 88]. As shown in 7(b), a turbo receiver operates by iteratively exchanging soft reliability information between the detector and the channel decoder, enabling both modules to refine their estimates progressively.
Unlike conventional receivers that perform detection and decoding in isolation, which is depicted in 7(a), a turbo receiver leverages the fact that the channel decoder can provide highly informative soft outputs, which in turn help the equaliser to suppress residual interference more effectively.
@@ 4.3.2 Turbo Receiver
4.4 Extension to MIMO-AFDM Systems
4.4.1 Channel Estimation in MIMO-AFDM
The principle of MIMO-AFDM transmission and channel estimation has been investigated in [102]. It can be interpreted as an extension of SISO-AFDM systems in Section 4-4.2. In a $N_{\mathrm{r}}\times N_{\mathrm{t}}$ system, it is possible to discriminate the channels if the pilots are distributed for any transmit antenna as
@@ 4.4.1 Channel Estimation in MIMO-AFDM
4.4.2 Special Cases of MIMO AFDM
Similarly to Section 4.2.3, we assume that the Doppler shift is small enough to be omitted, like in usual OFDM systems. Then, according to [103], it becomes possible to perform digital beamforming (precoding in DL or combining in UL) in the frequency domain. In fact, for any $n_{\mathrm{t}}=\{1,\ldots,N_{\mathrm{t}}\}$ and $n_{\mathrm{r}}=\{1,\ldots,N_{\mathrm{r}}\}$ we have
@@ 4.4.2 Special Cases of MIMO AFDM
In addition to the previous observation, it is worth mentioning that in the special cases of AFDM parameter [83] where $2N\lambda_{1}\in\mathbb{Z}$ and $\frac{1}{2\lambda_{1}}\in\mathbb{Z}$, subcarriers can be orthogonal in both the DAFT and the frequency domains. It is then possible to perform channel estimation in the frequency domain, as described in Section 4.2.3,along with frequency-domain beamforming.
4.5 Multiple Access and Coexistence with OFDM
@@
In the sequel, we briefly introduce affine frequency division multiple access (AFDMA) and its coexistence with orthogonal frequency division multiple access (OFDMA), AFDMA per block or per subcarrier, and MIMO-AFDMA.
@@
4.5.1 General AFDMA and Coexistence with OFDMA
@@
As an alternative, per-subcarrier AFDMA leverages specific AFDM parameter configurations where $2N\lambda_{1}\in\mathbb{Z}$ and $1/(2\lambda_{1})\in\mathbb{Z}$, as discussed in Section 4.2.3. Under these conditions, a data element mapped to an affine-frequency index $m$ is spread across specific frequency-domain indices $k$ satisfying $(k)_{2N\lambda_{1}}=(m)_{2N\lambda_{1}}$. By allocating subcarriers with index $(m)_{2N\lambda_{1}}=r$ to a given user, up to $K_{\mathrm{UE}}=2N\lambda_{1}$ users can be multiplexed. The resulting transmit signal is given by
@@ 4.5.2 MIMO-AFDMA
4.6 Phase Noise and Carrier Frequency Offset
@@ 4.6.1 Characterization of PHN and CFO
4.6.2 Waveform Resilience and Performance
@@
4.7 Degrees of Freedom via the Chirp Parameter Domain
@@ 4.7.1 Index Modulation
4.7.2 Physical Layer Security
@@
Table 2: Reusability Analysis: AFDM against the OFDM Legacy Infrastructure
@@ 4.7.3 PAPR Reduction
5 Reusability of AFDM with respect to OFDM
@@
A detailed breakdown of this compatibility is presented in Table 2. In the RF and Analog domain, the waveform’s statistical similarity to OFDM allows for the direct reuse of power amplifiers and transceivers without redesign. In the PHY Layer, while the core FFT hardware accelerators are preserved, the transition primarily necessitates software updates for channel estimation logic and phase rotation (chirping). Finally, compatible MAC integration is achieved through minimal extensions to existing structure and control signaling, validating AFDM as a high-performance yet low-risk evolutionary candidate for 6G+.
6 Conclusion

### v3

@@ Abstract
I Introduction
I-A Background
@@ I-B Why AFDM?
I-C Contributions of this article
@@
I-D Organization of this article
The remainder of this article is structured as follows: Section II introduces the system and channel models adopted throughout the article, including the novel formulation of a general FDFD channel input-output model based on physical pulse footprints to capture inter-sample coupling effects in practical transceiver implementations accurately; Section III details the transceiver architecture of AFDM, highlighting its compatibility and reusability with respect to OFDM, considering various aspects, such as prefixing, pulse shaping, and resource management; Section IV presents a variety of thorough analyses of AFDM in comparison with OFDM, including implementation complexity and scalability, channel estimation and detection architecture, robustness to impairments like phase noise (PHN) and CFO, and advanced functionalities including multiple-input multiple-output (MIMO), chirp-domain modulation and physical-layer security; Finally, Section V qualitatively summarizes the reusibility of the presented AFDM considerations and discussion across the RF, PHY, and higher layers, followed by concluding remarks in Section VI.
@@ II System Model
II-A Continuous-Time SISO Doubly Dispersive Channel
@@
II-B Discrete-Time Equivalent Model
@@
II-C Generalized Matrix-form Input-Output Relationship
@@
where $\phi_{\mathrm{cp}}(\cdot)$ is the waveform-dependent phase-offset function of the prefix samples, elaborated in Section III.
@@
II-D Extension to MIMO Channels
@@
Following the sampling procedure in Section II-B and the FDFD case, the discrete-time baseband channel between the $n_{\mathrm{t}}$-th transmit and $n_{\mathrm{r}}$-th receive antennas can be written as
@@
III On the Compatibility of AFDM over OFDM
@@
To that end, let us first introduce the input-output system model of the conventional OFDM and the AFDM modulator/demodulator, given the model of the discrete-time doubly dispersive channel in Section II-B.
III-A OFDM Signal Model
@@
Figure 3: Effective channel matrices $\mathbf{\Xi}$ for OFDM and AFDM, respectively on the first and second rows, in a 3-path doubly dispersive scenario with $N=64$. The physical channel parameters are defined by path coefficients $\{1,0.9,0.8\}$, normalized delays $\{1.3,3.25,5.96\}$, and normalized digital Dopplers $\{1.1,-2.3,0.85\}$, where for the IDID and IDFD cases, the delays and Dopplers are rounded to the nearest integers as relevant. For the AFDM implementation, chirp parameters are set to $c_{1}=(2(f_{\max}+1)+1)/2N$, with $f_{\max}=3$, and $c_{2}=1/2N$. It can be seen that on top of the discussed effects of the fraction taps in Section II-II-C, the different paths of the OFDM effective channel matrices are indistinguishable under Doppler shifts, leading to significant inter-carrier interference, while AFDM maintains a more distinguishable effective channel matrix, preserving better orthogonality.
@@
III-B AFDM Signal Model
@@
III-C MIMO Extension of OFDM and AFDM Signal Models
@@
III-D A Note on OFDM and AFDM Compatibility
@@
III-E Pulse-Shaping, Windowing, and Prefixing Blocks
@@
IV Analysis of the AFDM Transceiver over OFDM
IV-A Modulation Complexity and Scalability
@@
TABLE I: Computational complexity of key operations and total modulator/demodulator complexities in FLOPs, with $N=KL$.
@@
Comparing the modulator complexities including the pulse-shaping overheads of OFDM, AFDM, and the two OTFS implementations in (42) and (46), which are summarized in Table I, it is evident that both OTFS approaches introduce significantly higher computational burdens compared to the OFDM due to their multi-stage transform structures and 2D processing requirements under practical systems [^3] Regarding FFT dimension flexibility, OFDM and AFDM support arbitrary FFT/IFFT sizes $N$ with one-dimensional frequency-domain processing, enabling flexible numerology and resource allocation. In contrast, OTFS imposes strict integer divisibility constraints on the delay-Doppler grid dimensions $K,L\in\mathbb{N}$ such that $N=KL$, limiting design flexibility and potentially resulting in awkward FFT sizes for practical implementations and efficient computations., while AFDM maintains a modest overhead over OFDM and therefore stands out as a more computationally efficient alternative.
@@
IV-B Pilot Scheme and Channel Estimation
@@
IV-B1 Pilot Schemes
Due to the structure of the effective channel matrix $\mathbf{\Xi}^{\mathrm{AFDM}}$ as described in Section III-B, the usual pilot schemes for AFDM involve one or multiple pilot subcarriers [79, 29], each of which is isolated from other subcarriers by guard bands (in the DAFT domain) to capture the contribution of each path without interference properly.
@@ IV-B1 Pilot Schemes
IV-B2 Channel Estimation in the DAFT Domain
@@
$\mathbf{r}=\bigg(\sum_{p=1}^{P}\sum_{q=-B}^{B}h_{p}\cdot\mathbf{\Phi}_{p}\cdot\mathbf{V}^{f_{p}}\cdot g_{q}\cdot\mathbf{\Pi}^{\ell_{p}+q}\bigg)\cdot\mathbf{s}+\mathbf{w}.\vskip-2.15277pt$ (55)
@@
$\mathbf{r}=\Bigg(\sum_{u=1}^{P(2B+1)}\!\!\!h_{u}^{\prime}\cdot\mathbf{\Phi}_{u}\cdot\mathbf{V}^{f_{u}}\cdot\mathbf{\Pi}^{\ell_{u}^{\prime}}\Bigg)\cdot\mathbf{s}+\mathbf{w},\vskip-2.15277pt$ (56)
@@
IV-B3 Special Case: OFDM-like Estimation Under Low Doppler
@@
IV-C Signal Detection and Receiver Architectures
@@
IV-C1 One-shot Receivers (No outer iterations)
@@
IV-C2 Turbo Receiver
Alternative to the one-shot receivers where there are no outer iterations between the detector and the decoder, AFDM turbo receivers have been investigated in [100, 95, 88]. As shown in Fig. 7(b), a turbo receiver operates by iteratively exchanging soft reliability information between the detector and the channel decoder, enabling both modules to refine their estimates progressively.
Unlike conventional receivers that perform detection and decoding in isolation, which is depicted in Fig. 7(a), a turbo receiver leverages the fact that the channel decoder can provide highly informative soft outputs, which in turn help the equaliser to suppress residual interference more effectively.
@@ IV-C2 Turbo Receiver
IV-D Extension to MIMO-AFDM Systems
IV-D1 Channel Estimation in MIMO-AFDM
The principle of MIMO-AFDM transmission and channel estimation has been investigated in [102]. It can be interpreted as an extension of SISO-AFDM systems in Section IV-IV-B. In a $N_{\mathrm{r}}\times N_{\mathrm{t}}$ system, it is possible to discriminate the channels if the pilots are distributed for any transmit antenna as
@@ IV-D1 Channel Estimation in MIMO-AFDM
IV-D2 Special Cases of MIMO AFDM
Similarly to Section IV-B3, we assume that the Doppler shift is small enough to be omitted, like in usual OFDM systems. Then, according to [103], it becomes possible to perform digital beamforming (precoding in DL or combining in UL) in the frequency domain. In fact, for any $n_{\mathrm{t}}=\{1,\ldots,N_{\mathrm{t}}\}$ and $n_{\mathrm{r}}=\{1,\ldots,N_{\mathrm{r}}\}$ we have
@@ IV-D2 Special Cases of MIMO AFDM
In addition to the previous observation, it is worth mentioning that in the special cases of AFDM parameter [83] where $2N\lambda_{1}\in\mathbb{Z}$ and $\frac{1}{2\lambda_{1}}\in\mathbb{Z}$, subcarriers can be orthogonal in both the DAFT and the frequency domains. It is then possible to perform channel estimation in the frequency domain, as described in Section IV-B3,along with frequency-domain beamforming.
IV-E Multiple Access and Coexistence with OFDM
@@
In the sequel, we briefly introduce AFDMA and its coexistence with orthogonal frequency division multiple access (OFDMA), AFDMA per block or per subcarrier, and MIMO-AFDMA.
IV-E1 General AFDMA and Coexistence with OFDMA
@@
As an alternative, per-subcarrier AFDMA leverages specific AFDM parameter configurations where $2N\lambda_{1}\in\mathbb{Z}$ and $1/(2\lambda_{1})\in\mathbb{Z}$, as discussed in Section IV-B3. Under these conditions, a data element mapped to an affine-frequency index $m$ is spread across specific frequency-domain indices $k$ satisfying $(k)_{2N\lambda_{1}}=(m)_{2N\lambda_{1}}$. By allocating subcarriers with index $(m)_{2N\lambda_{1}}=r$ to a given user, up to $K_{\mathrm{UE}}=2N\lambda_{1}$ users can be multiplexed. The resulting transmit signal is given by
@@ IV-E2 MIMO-AFDMA
IV-F Phase Noise and Carrier Frequency Offset
@@ IV-F1 Characterization of PHN and CFO
IV-F2 Waveform Resilience and Performance
@@
IV-G Degrees of Freedom via the Chirp Parameter Domain
@@ IV-G1 Index Modulation
IV-G2 Physical Layer Security
@@
TABLE II: Reusability Analysis: AFDM against the OFDM Legacy Infrastructure
@@ IV-G3 PAPR Reduction
V Reusability of AFDM with respect to OFDM
@@
A detailed breakdown of this compatibility is presented in Table II. In the RF and Analog domain, the waveform’s statistical similarity to OFDM allows for the direct reuse of power amplifiers and transceivers without redesign. In the PHY Layer, while the core FFT hardware accelerators are preserved, the transition primarily necessitates software updates for channel estimation logic and phase rotation (chirping). Finally, compatible MAC integration is achieved through minimal extensions to existing structure and control signaling, validating AFDM as a high-performance yet low-risk evolutionary candidate for 6G+.
VI Conclusion